Sn = 1/1+√n + 1/2+√2n +.......+1/n+√n2


find limit n→∞ Sn

Grade:Upto college level

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Student

the sum can be written as ∑r=1r=n  1/(r + √rn)

= ∑r=1r=n  1/n(r/n + √r/n)

now replace 1/n by dx and r/n by x

limit of integral will be r/n

lower limit n tend to ∞ and r=1 will be 0

upper limit for r=n will be 1

it will convert into integral ∫01 dx/(x+√x) = ∫01dx/√x(1+√x)

now (1+√x)=t

and dx/2√x = dt

lower limit for x=0 , t=1 andupper limit for x=1, t=2

integral become ∫12 2dt/t= 2(lnt)12= 2ln2=ln4 answer

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..